# VoXtream: Full-Stream Text-to-Speech with Extremely Low Latency

### v1

@@ Abstract
We present VoXtream, a fully autoregressive, zero-shot streaming text-to-speech (TTS) system for real-time use that begins speaking from the first word. VoXtream directly maps incoming phonemes to audio tokens using a monotonic alignment scheme and a dynamic look-ahead that does not delay onset. Built around an incremental phoneme transformer, a temporal transformer predicting semantic and duration tokens, and a depth transformer producing acoustic tokens, VoXtream achieves, to our knowledge, the lowest initial delay among publicly available streaming TTS: 102 ms on GPU. Despite being trained on a mid-scale 9k-hour corpus, it matches or surpasses larger baselines on several metrics, while delivering competitive quality in both output- and full-streaming settings. Demo and code are available at https://herimor.github.io/voxtream.
@@ 1 Introduction
We introduce VoXtream, a fully autoregressive zero-shot TTS model that maps phoneme sequences to audio tokens with minimal delay. The architecture combines (1) an incremental Phoneme Transformer with dynamic phoneme look-ahead, (2) a Temporal Transformer (TT) that jointly predicts Mimi semantic and duration tokens, and (3) a Depth Transformer that generates acoustic tokens conditioned on outputs of TT and speaker embeddings. Unlike prior systems that require multi-stage decoding, complex alignment mechanisms, or large look-ahead windows, VoXtream combines these 3 components in a unified architecture. VoXtream begins outputting speech immediately after receiving the first word, achieving a first-packet latency as low as 102 ms on a modern GPU. Experiments conducted on SEED-TTS and LibriSpeech datasets show that the model generates highly intelligible speech, on par with state-of-the-art TTS models.
@@ 2 Method
Phoneme Transformer: The Phoneme Transformer (PT) is a decoder-only transformer that receives a phoneme sequence and encodes each phoneme with a corresponding embedding. The model operates incrementally, so its input grows with each new word from the text stream. For more natural prosody, PT is allowed to look ahead (LA) up to N phoneme tokens. A key advantage is dynamic LA: the model does not wait for N phonemes before generating speech tokens but starts immediately after the first word. The minimal LA value is determined by the text-buffer size, and the maximal value is limited to 10 phonemes. We convert the input text stream to phonemes using g2pE [^1] https://github.com/Kyubyong/g2p at the word level.
Temporal Transformer: The Temporal Transformer (TT) is an autoregressive transformer conditioned on the audio tokens and the corresponding phoneme sequence. Audio tokens are extracted by the Mimi [19] codec at 12.5 Hz. Due to the low frame rate, we assign up to two phonemes to each audio frame. Alignment between the acoustic prompt and the phoneme sequence is obtained using the Montreal Forced Aligner (MFA) [20]. Following [19], we apply an acoustic delay of one step for stability. The TT outputs the first codebook of Mimi (semantic tokens) and a duration token. Inspired by [21], we adopt monotonic phoneme alignment, but instead of predicting the specific phoneme, we predict a duration token that encodes a shift flag (stay/go, the first number in duration token), where “stay” means that we continue generating of current phoneme in the next frame, or switch to the next phoneme otherwise, and the number of phonemes (1 or 2, the second number in duration token), corresponding to slower or faster pronunciation. Semantic and duration tokens are predicted by a single classification head and sampled from the joint distribution. The index of acoustic tokens corresponds to the temporal and codebook positions.
@@
The Mimi decoder converts semantic and acoustic tokens for each frame into 80 ms of speech in a streaming fashion. We train VoXtream by minimizing the negative log-likelihood of TT and DT outputs.
@@ 3.1 Experimental Setup
VoXtream:Out Phone 9k EN 471M 3.82 0.529 3.88 3.09 0.461 4.08 53.4 $\pm$ 2.5
VoXtream:Full Phone 9k EN 471M 3.81 0.529 3.90 3.15 0.458 4.07 51.9 $\pm$ 2.6
@@
Dataset: Our training corpus is based on the Emilia [23] and HiFiTTS-2 [24] datasets, covering both spontaneous and read styles. We selected 4.5k hours from each, totaling 9k hours to be comparable to [25] in dataset size. For Emilia, we applied additional diarization to remove multi-speaker utterances and discarded utterances with invalid automatic transcripts. We also used NISQA [26] to remove low-quality utterances. Phoneme alignments were obtained with MFA [20]. Speech tokenization used the streaming Mimi [19] codec at 24 kHz.
@@
Baseline Models: We chose multiple AR models with publicly available weights for comparison. For the non-streaming scenario we selected CosyVoice [2], Spark-TTS [3], Llasa [4], VoiceStar [29], VoiceCraft [25], XTTS [5], CosyVoice2 [16], and FireRedTTS-1S [6]. We split them into Large and Mid groups by training data size for fair comparison.
@@ 3.2 Evaluation
To evaluate naturalness, we conducted two user studies on the Prolific platform. Each study included multiple attention checks, and participants were compensated at an average rate of 9 GBP per hour. For a fair comparison, we sampled 100 unique sentences on which all systems achieved 0% WER. The first study focused on short-form TTS. We recruited 40 native listeners and asked them to rate the naturalness of each system on a 0–100 scale, following a MUSHRA-like protocol. The second study focused on full-stream evaluation. Here, 30 native listeners were asked to select their preferred system or indicate “no preference”.
@@
Table 3: Performance in FP16 on an A100 GPU. TC denotes torch.compile and DS denotes DeepSpeed. Metrics for XTTS-v2 are for output streaming; other models are full-stream.
@@ 6 Acknowledgements
This work was partially supported by the Wallenberg AI, Autonomous Systems and Software Program (WASP) funded by the Knut and Alice Wallenberg Foundation.

### v2

@@ Abstract
We present VoXtream, a fully autoregressive, zero-shot streaming text-to-speech (TTS) system for real-time use that begins speaking from the first word. VoXtream directly maps incoming phonemes to audio tokens using a monotonic alignment scheme and a limited look-ahead that does not delay onset. Built around an incremental phoneme transformer, a temporal transformer predicting semantic and duration tokens, and a depth transformer producing acoustic tokens, VoXtream achieves, to our knowledge, the lowest initial delay among publicly available streaming TTS: 102 ms on GPU. Despite being trained on a mid-scale 9k-hour corpus, it matches or surpasses larger baselines on several metrics, while delivering competitive quality in both output- and full-streaming settings. Demo and code are available at https://herimor.github.io/voxtream.
@@ 1 Introduction
We introduce VoXtream, a fully autoregressive zero-shot TTS model that maps phoneme sequences to audio tokens with minimal delay. The architecture combines (1) an incremental Phoneme Transformer with limited phoneme look-ahead, (2) a Temporal Transformer (TT) that jointly predicts Mimi semantic and duration tokens, and (3) a Depth Transformer that generates acoustic tokens conditioned on outputs of TT and speaker embeddings. Unlike prior systems, the proposed method encodes text incrementally, making it possible to generate speech with a small look-ahead, as past text states are updated as more context is available. VoXtream begins outputting speech immediately after receiving the first word, achieving a first-packet latency as low as 102 ms on a modern GPU. Experiments conducted on SEED-TTS and LibriSpeech datasets show that the model generates highly intelligible speech, on par with state-of-the-art TTS models.
@@ 2 Method
Phoneme Transformer: The Phoneme Transformer (PT) is a decoder-only transformer that receives a phoneme sequence and encodes each phoneme with a corresponding embedding. The model operates incrementally, so its input grows with each new word from the text stream. For more natural prosody, PT is allowed to look ahead (LA) up to N phoneme tokens. A key advantage is limited LA, where the model does not wait for N phonemes before generating speech tokens but starts immediately after the first word. The minimal LA value is determined by the text-buffer size, and the maximal value is limited to 10 phonemes. We convert the input text stream to phonemes using g2pE [^1] https://github.com/Kyubyong/g2p at the word level.
Temporal Transformer: The Temporal Transformer (TT) is an autoregressive transformer conditioned on the audio tokens and the corresponding phoneme sequence. Audio tokens are extracted by the Mimi [8] codec at 12.5 Hz. Due to the low frame rate, we assign up to two phonemes to each audio frame. Alignment between the acoustic prompt and the phoneme sequence is obtained using the Montreal Forced Aligner (MFA) [22]. Following [8], we apply an acoustic delay of one step for stability. The TT outputs the first codebook of Mimi (semantic tokens) and a duration token. Inspired by [15], we adopt monotonic phoneme alignment, but instead of predicting the specific phoneme, we predict a duration token that encodes a shift flag (stay/go, the first number in duration token), where “stay" means that we continue generating of current phoneme in the next frame, or switch to the next phoneme otherwise, and the number of phonemes (1 or 2, the second number in duration token), corresponding to slower or faster pronunciation. Semantic and duration tokens are predicted by a single classification head and sampled from the joint distribution. The index of acoustic tokens corresponds to the temporal and codebook positions.
@@
The Mimi decoder converts semantic and acoustic tokens for each frame into 80 ms of speech in a streaming fashion. We train VoXtream by minimizing the negative log-likelihood of TT and DT outputs. An external aligner is used to obtain the initial alignment between the acoustic and text prompts.
@@ 3.1 Experimental Setup
VoXtream:Out Phone 9k EN 441M 3.82 0.529 3.88 3.09 0.461 4.08 53.4 $\pm$ 2.5
VoXtream:Full Phone 9k EN 441M 3.81 0.529 3.90 3.15 0.458 4.07 51.9 $\pm$ 2.6
@@
Dataset: Our training corpus is based on the Emilia [16] and HiFiTTS-2 [19] datasets, covering both spontaneous and read styles. We selected 4.5k hours from each, totaling 9k hours, to be comparable to [27] in dataset size. For Emilia, we applied additional diarization to remove multi-speaker utterances and discarded utterances with invalid automatic transcripts. We also used NISQA [24] to remove low-quality utterances. We used the CMU dictionary to map text into phonemes and aligned them to audio with MFA [22]. Speech tokenization used the streaming Mimi [8] codec at 24 kHz.
@@
Baseline Models: We chose multiple AR models with publicly available weights for comparison. For the non-streaming scenario we selected CosyVoice [11], Spark-TTS [31], Llasa [35], VoiceStar [26], VoiceCraft [27], XTTS [3], CosyVoice2 [12], and FireRedTTS-1S [14]. We split them into Large/Mid groups by training data size for fair comparison.
@@ 3.2 Evaluation
To evaluate naturalness, we conducted two user studies on the Prolific platform. Each study included multiple attention checks, and participants were compensated at an average rate of 9 GBP per hour. For a fair comparison, we sampled 100 unique sentences on which all systems achieved 0% WER. The first study focused on short-form TTS. We recruited 40 native listeners and asked them to rate the naturalness of each system on a 0–100 scale, following a MUSHRA-like protocol. The second study focused on full-stream evaluation. Here, 30 native listeners were asked to select their preferred system or indicate “no preference".
@@
Table 3: Performance in FP16 on an A100 GPU. TC denotes torch.compile and DS denotes DeepSpeed. XTTS-v2 metrics are for output streaming; other models are full-stream.
@@ 6 Acknowledgements
This work was partially supported by the Wallenberg AI, Autonomous Systems and Software Program (WASP) funded by the Knut and Alice Wallenberg Foundation (KAW) and by the Industrial Strategic Technology Development Program (grant no. 20023495) funded by MOTIE, South Korea. Computations were enabled by the supercomputing resource Berzelius provided by KAW and NSC at Linköping University.